OpenStudy (anonymous):

use the definition of a linear operator to show that the function T:R^2 -> R^2 given the formula T(x,y)= (x+2y , 3x-y) is a linear operator.

OpenStudy (anonymous):

pls show step by step. esp for that one of T(u+v) = T(u) + T(v)

OpenStudy (anonymous):

first of all check ur funtion u have T(x,y) =(x+2y,3x-y) ur let \[u=(x _{1},y_{1}) and v=(x_{2},y _{2})\] from here T(U+V)= \[T(x _{1}+x _{2},y _{1}+y _{2})\] if u r substituting u don't put T by substituting i mean where u see x with subscript 1 or 2 u put (x+2y) with the corresponding subscript where u see y with subscript 1or 2 u put (3x-y) with the corresponding subscript now u'll have \[(x _{1}+2y _{1}+x _{2}+2y _{2};3x _{1}-y _{1}+3x_{2}+y _{2})\] separate them \[(x _{1}-2y _{1};3x _{1}-y _{1}) + (x _{2}-2y _{2};3x _{2}-y _{2})\] check this interesting part \[(x _{1}+2y _{1};3x _{1}-y _{1}) =T(x _{1};y _{1})\] but \[(x _{1};y _{1})=u\] then u'll have T(u) similarly do the same technique to v

OpenStudy (anonymous):

the part wher ur separating.. what defn did u use to distribute + over ;